An ideal Otto engine, operating on the hot-air standard with k=1.34, has a

compression ratio of 5. At the beginning of compression the volume is 6ft3

, the

pressure is 15 psia and temperature is 100F. during constant – volume heating , 500

Btu are added per cycle. Compute Wnet, Qr, thermal Efficiency, and mean

effective pressure.


1
Expert's answer
2022-03-10T18:01:11-0500

Solution;

Given;

k=1.34

"V_1=6ft^3"

"P_1=15psia"

"r=5"

"T_1=100F=560\u00b0R"

"Q_a=500Btu"

From the relation;

"\frac{T_2}{T_1}=(\frac{V_1}{V_2})^{k-1}=r^{k-1}"

"T_2=560\u00d75^{0.34}=968\u00b0R"

The mass can be obtained as;

"m=\frac{PV}{RT}"

"m=\frac{15×144×6}{53.34×560}=0.4339lb"

We can find Cv as;

"C_v=\frac{R}{k-1}=\frac{53.34}{0.34}=156.88ft/lb°R"

Heat added;

"Q_a=mC_v\Delta T"

"Q_a=mC_v(T_3-T_2)"

From which we obtain "T_3" as;

"T_3=\frac{Q_a}{mC_v}+T_2"

"T_3=\frac{500×778}{0.4339×156.88}+968"

"T_3=6682.69\u00b0R"

"P_2=P_1r^k"

"P_2=15\u00d75^{1.34}=129.63psia"

"P_3=P_2(\frac {T_3}{T_2})=129.63(\frac{6682.69}{968})"

"P_3=894.93psia"

"T_4=T_3(\frac {1}{r})^{k-1}"

"T_4=6682.69(\frac15)^{0.34}"

"T_4=3866.47\u00b0R"

"Q_r=mC_v(T_1-T_4)"

"Q_r=0.4339\u00d7156.88(560-6682.69)"

"Q_r=535.70Btu"

Efficiency;

"\eta=1-\frac{1}{r^{k-1}}"

"\eta=1-\frac{1}{5^{0.34}}=0.4214"

"W_{net}=\eta×Q_a"

"W_{net}=0.4214\u00d7500=210Btu"

Mean effective pressure;

"P_m=\frac{W_net}{V_1-V_2}"

"V_2=\frac{V_1}{r}=\frac{6}{5}=1.2ft^3"

"P_m=\frac{210×778}{(6-1.2)×144}"

"P_m=236.37psia"